OpenStudy (anonymous):

The number of milligrams of a diabetic drug in a person's system after t hours is given by the function D(t)=20e^(-0.4t). How much of the drug is in the person's system after 5 hours?

OpenStudy (anonymous):

@jim_thompson5910 @satellite73 @Mertsj

jimthompson5910 (jim_thompson5910):

Plug in t = 5 and evaluate. D(t)=20e^(-0.4t) D(5)=20e^(-0.4*5) D(5)=20e^(-2) D(5)=20(0.13533528) D(5)=2.7067056 So roughly 2.7067056 mg of the drug is left in the system.

OpenStudy (anonymous):

Thank you soooo much! =)

jimthompson5910 (jim_thompson5910):

sure thing :)